Evaluvate integral 0 to 1 tan inverse xdx.
1
Expert's answer
2020-05-11T19:18:25-0400

Here, the objective is evaluate the definite integral using integration by parts.


Let "u=\arctan(x),dv=dx" , then


"du=\frac{1}{1+x^2}dx"


"v=x"


So, use the integration by parts as,


"\int_{0}^1(\arctan(x))dx=uv-\int(v)du"


"=[x\arctan(x)]_{0}^1-\int_{0}^1\frac{x}{1+x^2}dx"


"=\frac{\pi}{4}-\frac{1}{2}\int_{0}^1\frac{2x}{1+x^2}dx"


Here, "g'(x)=2x," if "g(x)=1+x^2" , so using fundamental theorem of calculus, the integral is,


"=\frac{\pi}{4}-\frac{1}{2}[ln|1+x^2|]_{0}^1"


"=\frac{\pi}{4}-\frac{1}{2}[ln|1+1|-ln|1|]"


"=\frac{\pi}{4}-\frac{1}{2}[ln2-0|]"


"=\frac{\pi}{4}-\frac{1}{2}ln2"


Therefore, the value of the definite integral is "\int_{0}^1(\arctan(x))dx=\frac{\pi}{4}-\frac{1}{2}ln2"